OpenStudy (pretzel2000):

Braxton and Maggie babysit children for extra money over the summer. Braxton's income is determined by f(x) = 9x + 10, where x is the number of hours. Maggie's income is g(x) = 6x + 15. If Braxton and Maggie were to combine their efforts, their income would be h(x) = f(x) + g(x). Assume Braxton works 5 hours. Create the function h(x), and indicate if Braxton will make more money working alone or by teaming with Maggie.

OpenStudy (legomyego180):

Any ideas?

OpenStudy (pretzel2000):

Its a multiple choice question but I have no clue how to solve it h(x) = 15x + 25, team with Maggie h(x) = 3x + 5, work alone h(x) = 3x + 5, team with Maggie h(x) = 15x + 25, work alone

OpenStudy (legomyego180):

Braxton's income: \[f(x)=9x+10\] Maggies income: \[g(x)=6x+15\] Combined income: \[h(x) = f(x) + g(x) = (9x+10) + (6x+15)\] Can you simplify the combined income?

OpenStudy (pretzel2000):

so if I combine the income the answer would be A) h(x)=15x+25?

OpenStudy (legomyego180):

Thats the equation but there are two answers with that equation, need to find out which way will earn braxton the most cash

OpenStudy (legomyego180):

So alone Braxton will make \[9(5)+10=$55\] Together with whats her face you'll plug 5 into h(x) but you have to split the payment 50% with emily so we will divide it by two. \[\frac{ (9(5)+10)+ (6(5)+15) }{ 2 }\]